### v2

# Self-assembling clusters of particles on a shrinking liquid surface

Abstract
After rainfall, pine needles are observed to float on the surface of small puddles. As the water evaporates, these pine needles exhibit spontaneous self-assembly into distinct clusters. Motivated by this natural phenomenon, we conducted experimental investigations into the dynamic evolution of particles during the shrinkage of a liquid surface. Our experiments revealed that particles tend to aggregate, forming distinct clusters as the liquid boundary shrinks. We systematically examined the emergence of these clusters and explored how their sizes and numbers evolve with changes in packing fraction. Furthermore, we analyze particle rotation during the process and discuss the formation of the final configuration comprising clusters of various orientations. Complementary numerical simulations demonstrate qualitative agreement with our experimental findings. Our study contributes to a better understanding of self-assembly phenomena involving macroscopic particles on shrinking liquid surfaces.
preprint: APS/123-QED
I INTRODUCTION
It is of significant interest to understand the ability of particles to self-assemble into complex structures. By exploring how this collective behavior emerges, we can gain insight into fundamental questions of how matter organizes itself, essential to condensed matter physics, materials science, and even our understanding of life itself Manoharan (2015). Self-assembly of particles is mostly observed in colloidal systems (e.g., Refs. Boles et al. (2016); Li et al. (2016, 2021); Peng et al. (2023)). Recent advancements in synthesis have led to the production of a diverse array of anisotropic particles, which readily organize into various superlattices under the appropriate experimental conditions (e.g., Refs. Henzie et al. (2012); Young et al. (2013); Gong et al. (2017); Forster et al. (2011); Hosein et al. (2010); Meijer et al. (2017); Miszta et al. (2011)). When the predominant interparticle interaction is the hard core potential of particle shape and other interactions are weak, colloidal particles can be approximated as hard particles. Simulations of hard colloids (e.g., Refs. Torquato and Jiao (2009); Marechal and Dijkstra (2010); Gang and Zhang (2011); Agarwal and Escobedo (2011); Smallenburg et al. (2012); Avendaño and Escobedo (2012); Gantapara et al. (2013); Bernard and Krauth (2011); Wan and Glotzer (2018); Wan (2022); Liu et al. (2023)) predict complex crystals from an even larger variety of anisotropic shapes.
As particles increase in size, however, they become less susceptible to motion induced by thermal fluctuations of the solvent molecules, limiting self-assembly in everyday scenarios. To the best of our knowledge, the only exception where the self-assembly of macroscopic particles is due to the presence of the liquid is the Cheerios effect, which refers to the phenomenon of floating objects appearing to attract one another (e.g., Refs. Nicolson (1949); Vella and Mahadevan (2005); Grzybowski et al. (2009); Vella (2015); Protière (2023); Bowden et al. (1997)). This effect is observed, for example, when pieces of breakfast cereal floating on the surface of a bowl tend to clump together, hence the name. In this case, the weight of the particles causes them to deform the air-milk interface. To minimize the overall distortion of the surface, surface tension forces the particles together.
After rainfall it is frequently observed that fallen pine needles in a small puddle first float on the water surface. As water evaporates from the puddle, however, pine needles spontaneously form distinct clusters of varying orientations [see Fig. 1(a)]. The dashed blue line in the figure indicates the outline of a cluster. In this cluster, pine needles have a similar orientation indicated by the blue arrow. Here we ask the question: how do these clusters of various orientations form and settle into well-defined structures?
We attribute this clustering phenomenon to two main factors: the Cheerios effect, which brings the needles together, and the evaporation of water, leading to a reduction in surface area and a consequent increase in packing density. To systematically investigate this process, wherein macroscopic particles at the liquid surface self-assemble while the surface area reduces, we conducted experimental studies [see Fig. 1(b)]. We characterized the emergence of clusters, examined the evolution of cluster sizes and their number as a function of packing fraction, and analyzed particle rotation during the process. Additionally, we conducted numerical simulations that account for the adhesion of particles and the reduction of surface area, with results qualitatively agreeing with our experimental findings. We discuss the formation of the final configuration comprising clusters of various orientations.
Figure 1: (a) Pine needles in a small puddle after rain, exhibiting self-assembly into clusters with particles in a cluster having similar orientations. The blue dashed line outlines a cluster, and the blue arrow indicates their orientation. (b) A snapshot of particles in our experiment at a relatively high density. Similar to (a), the blue dashed line outlines a cluster, and the blue arrow shows their orientation. (c) Photograph of the experimental setup. (d) Illustration depicting the experimental setup.
II Experimental Setup
Figure 1(c) shows the experimental setup which consists of two identical buckets arranged vertically. The lower bucket rests on the floor while the higher one sits atop a shelf to generate a vertical gradient. The elevated bucket serves as a water-filled reservoir. The two buckets are connected by a pipe equipped with a valve and so water siphons from the higher bucket to the lower one. In the lower bucket a metal cone in the form of a frustum is strategically positioned on three small blocks, elevating it above the bucket’s bottom [see Fig. 1(d) for an illustration of the experimental setup]. This setup enables water to enter from the base, connecting the cone with the bucket. Initially, we fill the lower bucket with water until it submerges the three supporting blocks of the metal cone. To confine particles on the liquid surface in the subsequent step, we add a layer of corn oil to the water surface. The oil, with a density of 0.91 g/cm³, forms a layer approximately 1.57 cm thick. Subsequently, we place 150 white polylactic acid (PLA) particles sparsely across the oil surface. These PLA particles, with dimensions of 35 mm $\times$ 7 mm $\times$ 7 mm, are partially filled inside (see section A of the Materials and Methods for particle details), resulting in a density of 0.62 g/cm³, allowing them to float on the oil surface. Additionally, the four major faces of the particles are marked with red markers, enabling us to track their positions and orientations through image analysis. A camera is positioned overhead to record the experiments. The initial diameter of the oil surface within the cone measures approximately 44 cm, resulting in an initial packing density of about 0.24 [Fig. 2(a)]. The experiment starts with opening the valve, enabling water to flow from the higher bucket to the lower one at the slowest possible rate. As the surface level rises, its area decreases due to the cone’s shape.
III Results and discussion
Figure 2: (a)-(e) Sequential photographs capturing the system’s evolution at 0, 13, 59, 75, and 93 minutes, respectively, showcasing the merging of nearby particles or clusters into larger entities, highlighted by red, yellow, and blue windows. (f)-(j) Visualization of cluster identification in the system, with cluster colors representing their respective sizes (number of particles). The maroon, palevioletred, and orchid windows in (j) denote the three largest clusters, each containing 8 particles. (k)-(o) Corresponding configurations in simulations at matching packing densities to (f)-(j). In (k), maroon, palevioletred, and orchid particles represent those forming the largest clusters in (j), while others are depicted in grey. (p)-(t) Distribution of cluster size $n_{c}$ corresponding to the configurations depicted in (f)-(j) for experiments and (k)-(o) for simulations.
Figure 3: (a)-(d) Distribution of rotation angle $\theta$ obtained from experimental data, corresponding to the configurations illustrated in Fig. 2(b)-2(e), respectively. Each bin represents a 10-degree width. (e)-(h) Distribution of rotation angle $\theta$, extracted from simulations conducted at identical packing densities as in (a)-(d), corresponding to configurations depicted in Fig. 2(l)-2(o), respectively.
Figure 4: (a) Average cluster size $<n_{c}>$ as a function of packing density $\phi$. (b) Average rotation angle $<\theta>$ as a function of packing density $\phi$. (c) Illustration of cluster formation with various orientations. Colors indicate orientation.
Figures 2(a)-2(e) depict the evolution of particle configurations as the surface level rises. The packing density collectively increases as the surface area decreases, with neighboring particles adhering to form clusters. We highlight several examples using red, yellow, and blue windows in the figures. To gain deeper insight into cluster formation, we analyze experimental configurations (see section B of the Materials and Methods for image analysis) and identify clusters within the configurations (see section E of the Materials and Methods for cluster identification). Clusters are color-coded based on their size, with particles in the same cluster sharing the same color, as illustrated in Figs. 2(f)-2(j). Initially, the configuration displays only one largest cluster of size 6 [highlighted in yellow in Fig. 2(f)]. As the experiment progresses, larger clusters emerge. The experiment lasts for approximately 93 minutes. The final configuration reaches a packing density of approximately 57%, featuring three large clusters of size 8 [Fig. 2(j)]. These three largest clusters are marked in red and emphasized by maroon, palevioletred, and orchid windows, respectively.
To further elucidate the system’s evolution, we perform Monte Carlo simulations to investigate the self-assembly process. Here, we model particles as two-dimensional hard rectangles and employ hard particle Monte Carlo simulations (e.g., Refs. Haji-Akbari et al. (2009); Damasceno et al. (2012); Ni et al. (2012); Anderson et al. (2017); Wan and Glotzer (2021); Qin et al. (2023); Lei et al. (2018); Klotsa et al. (2018); Wan et al. (2019)) augmented with an adhesion mechanism between particles inspired by the Cheerios effect. In our simulations, when the outermost particles of two adjacent clusters (or individual particles) are within an approximate interaction range dictated by the Cheerios effect Vella and Mahadevan (2005), and they exhibit similar orientation, we allow them to adhere with a specified probability. For simplicity, we only consider the attraction between the two outermost particles and focus solely on the interaction between their elongated sides, neglecting influences from other particles or clusters. The final overlap length of the particles, as a function of the initial overlap length, is determined from experiments on two-particle attraction. To accommodate for area shrinkage the circular boundary was gradually contracted in the simulations. We execute the hard particle Monte Carlo simulations with the adhesion mechanisms until finding configurations that accommodate the boundary contraction. These configurations are recorded at specific packing densities. Detailed simulation procedures are provided in Section D of the Materials and Methods.
To align closely with the experimental setup, we initialized the simulation with the same configuration as in the experiment, depicted in Fig. 2(k), mirroring Fig. 2(f). In Fig. 2(k), we illustrate particles forming the three largest clusters in the final experimental configuration [Fig. 2(j)] in maroon, palevioletred, and orchid, respectively, reflecting the color scheme of the corresponding windows in Fig. 2(j), while other particles are depicted in grey. Notably, particles forming the same clusters in the final configuration originate from close proximity in the initial configuration. Figures 2(l)-2(o) display the system’s configuration in simulations at packing densities matching experiment. It is evident from the figures that as the packing density increases the system tends to include more clusters of larger sizes, qualitatively aligning with experimental observations. Figures 2(p)-2(t) illustrate the distribution of cluster size $n_{c}$ (the number of particles in a cluster) in both experiments and simulations. The distributions from both experiments and simulations reveal that with increasing packing density, the number of small clusters decreases, while the number of larger clusters gradually increases.
We proceed to analyze particle rotation during the self-assembly process. We assign each particle a pointing direction at the beginning (0 min) and track its direction throughout the evolution. Figures 3(a)-3(d) plot the distribution of rotation angle $\theta$ at times as shown in Figs. 2(b)-2(e). Here, we only consider the magnitude of the rotation, but not its direction. As can be seen from the figure, over the evolution, particles with larger rotation angles emerge. However, the rotation remains relatively small. Notably, except for one particle with $\theta$ slightly above 90 degrees in Fig. 3(a) and reaching 350 to 360 degrees in Figs. 3(c) and 3(d), all other particles exhibit rotation angles below 90 degrees in Figs. 3(a)-3(d). Figures 3(e)-3(h) plot $\theta$ obtained from simulation at the matching densities as in 3(a)-3(d). We can see that simulation results exhibit a similar trend.
We show the statistics of the average cluster size $<n_{c}>$ and the average rotation angle $<\theta>$ as a function of packing density $\phi$ in Fig. 4. As can be seen from Fig. 4(a), the initial average number of particles in a cluster is approximately 1.3; the final value in experiments goes up to about 2.4. The average number of particles in a cluster in both experiments and simulations demonstrates qualitative agreement. Regarding the average rotation angle, the average rotation angle from the experiment has a significant jump at packing density $\phi=0.27$ and shows a value of about 17.4 degrees (red circle); after that, the increasing trend decreases, with $<\theta>$ for the final configuration at about 27.6 degrees. The simulation results show a similar trend. If we exclude the particle that has a super-large rotation angle and then calculate the average, the experimental results (green circle) and the simulation results are closer. This observation, that particles have a relatively small average rotation angle, alongside the tendency for clusters to form from nearby particles, suggests cluster growth occurs through the absorption of similarly oriented particles or small clusters, with minimal movement or rotation. When particles or clusters adhere together, they show more similar orientations. However, as packing density increases, misaligned clusters unable to adhere tend to stack, resulting in the final configuration comprising clusters of varied orientations [Fig. 4(c)].
The system we have studied fundamentally differs from the self-assembly of colloidal particles in suspension subjected to thermal noise. For example, hard rod-like colloidal particles exhibit rapid translational and rotational Brownian motions in the isotropic phase, forming small, transient side-by-side clusters that quickly collapse due to thermal agitation or collisions with surrounding particles Maeda and Maeda (2003). In our system, in contrast, once two particles, or a particle and a cluster, adhere, the bonds remain intact. We rarely observe changes in overlap length, except at relatively high packing densities where particles become more compressed. Collisions between clusters may increase the overlap length. This stable bonding is attributed to the strong interaction at contact Vella and Mahadevan (2005), which stabilizes the adhered structure and causes it to move as a single entity. Furthermore, the interaction decays rapidly at distances much greater than the capillary length, which in our case is several millimeters. Consequently, as the surface area shrinks, the distance between particles decreases, enhancing the Cheerios effect and promoting the aggregation of clusters.
IV Conclusions
In conclusion, inspired by the observation of fallen pine needles self-assembling into clusters in small puddles after rainfall, we conduct experimental investigations into the dynamic evolution of particles as the liquid surface undergoes shrinkage. We observe that particles tend to adhere to one another, forming distinct clusters as the surface area reduces. We identify the emergence of clusters and systematically explore how cluster sizes and the number of clusters evolve as a function of packing fraction. Furthermore, we analyze particle rotation during the process and discuss the formation of the final configuration comprising clusters of various orientations. Complementary numerical simulations that capture the two main factors, the adhesion between particles and the reduction of surface area, demonstrate qualitative agreement with our experimental findings. Our study contributes to a better understanding of the self-assembly phenomena occurring when macroscopic particles interact on a liquid surface whose area is shrinking.
V Materials and Methods
V.1 Preparation of particles
White PLA particles were produced via 3D printing using a Flashforge Guider II S model. Each particle measures 35 mm in length, 7 mm in width, and 7 mm in height, featuring a centrally located groove measuring 30 mm in length, 3 mm in width, and 1 mm in depth along each of its four major sides. The particles were deliberately designed to be hollow, resulting in a filling ratio of 15%, thereby improving their buoyancy on liquids. The detailed structure of the particle is illustrated in Fig. 5. Please note that the colors depicted in Fig. 5 do not represent the actual colors of the particles; the real particles are white, with grooves distinctly marked in red using a red marker to enhance visibility for the following image analysis.
Figure 5: Structure of a particle. (a) External structure. Each face of the particle exhibits a central groove. Solid fills are denoted in orange, inner walls in green, and outer walls in flesh color. (b) Internal structure. The particle’s hollow interior reveals a hexagonal filling shape (highlighted in yellow). The cross-sectional view is presented on the right. Colors maintain consistent color coding as in (a).
V.2 Image analysis
We utilized OpenCV (Open Source Computer Vision Library, implemented in Python) for particle identification. This software enables the detection of the minimum rectangle encapsulating the red rectangular marker within the field of view. After obtaining the coordinates of the four vertices of the red markers, we ascertain both the center positions and orientations of the particles.
V.3 Two-particle attraction
To investigate the attraction between two particles, we align them parallel to each other with their large faces facing each other. Maintaining a horizontal separation of $d=1.2$ cm, we systematically vary the initial vertical overlap length $x$ from 1/7 to 6/7 of the particle length, in increments of 1/7 as equal divisions. Upon releasing the particles, they adhere to each other, and we measure the final overlap length $y$ for each initial overlap length $x$, as depicted in Fig. 6(a). The black line represents the linear fit of the first five experimental data points, where the final overlap $y$ values are less than the particle length. The blue line denotes the particle length.
Figure 6: (a) Final overlap length $y$ as a function of initial overlap length $x$ when two particles are released at a fixed horizontal distance $d=1.2$ cm. The inset illustrates the initial and final states. The black line represents the linear fit of the experimental data based on the first five data points, while the blue line denotes the particle length. (b) Determination of the adhesion of two clusters in simulations. When the outermost particles (marked in blue) of two adjacent clusters form an angle $\alpha\leq 5^{\circ}$, we rotate each cluster by $\alpha/2$ to align them parallelly. If the horizontal distance $d\leq 1.2$ cm and the vertical overlap length $x$ is greater than 1/7 of the particle length, the two clusters are allowed to adhere with a probability of $p=0.02$. The final vertical overlap length $y$ is determined by the black and blue lines in (a).
V.4 Numerical simulations
We employ Monte Carlo simulations to investigate the self-assembly of two-dimensional PLA particles as the overall area shrinks. These simulations adhere to the standard hard particle Monte Carlo approach, with the inclusion of an adhesion mechanism to replicate attractive interactions, commonly known as the Cheerios effect Vella and Mahadevan (2005); Protière (2023). The particles are modeled as two-dimensional rectangles, and each Monte Carlo step involves $N$ trial moves, encompassing translations and rotations of the particles Haji-Akbari et al. (2009); Wan and Glotzer (2021). We set the length of the long side of the particle as the unit length in simulations. Translation steps are fixed at 0.063, approximately 2.2 mm, while rotation steps are 0.033 radians, roughly 1.9 degrees. Overlap checks between particles are conducted using the hard particle Monte Carlo module in HOOMD-blue Anderson et al. (2016), and moves resulting in overlap or boundary intersection are rejected.
To account for the attraction between particles, we incorporate an adhesion mechanism. For simplicity, we consider adhesion only between the long sides of particles and attraction between two particles. Specifically, when the outermost particles of two adjacent clusters (or individual particles) form an angle of $\alpha\leq 5^{\circ}$, we rotate each cluster by $\alpha/2$ to align them in parallel [see an illustration in Fig. 6(b)]. If the horizontal distance $d\leq 1.2$ cm and the vertical overlap length $x$ exceeds 5 mm (1/7 of the particle length), we allow the two clusters to adhere to each other with a probability of $p=0.02$. The final vertical overlap length $y$ is determined by the black line fitting in Fig. 6(a) if $y$ is less than the particle length; otherwise, $y$ is set to the particle length, as indicated by the blue line in the figure.
We initiate simulations with $N=150$ rectangular particles, maintaining the same aspect ratio as experimental conditions, within a circular boundary. To closely mimic the experimental setup, we initialize the system with the same configuration as in the experiment. To accommodate liquid surface shrinkage, the circular boundary is rescaled to 0.995 times its initial value. If the rescaled boundary overlaps with particles, rescaling is rejected, and 10 Monte Carlo steps with the adhesion mechanism are performed. This process is repeated until successful boundary rescaling. Configurations are recorded at specific packing densities.
V.5 Identification of clusters
We identify particles belonging to the same cluster in both experiments and simulations based on the following criteria: (1) The orientation difference between two particles is less than 0.4 radians; (2) The distance between the mass centers of the two particles is less than 0.89 units; (3) The horizontal distance and vertical overlap length after aligning the two particles parallel [denoted as $d$ and $x$ in Fig. 6(b)] are less than 0.268 units and greater than 0.37 units, respectively.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant No. 12274330) and the Knowledge Innovation Program of Wuhan-Shuguang (Grant No. 2022010801020125). D.W. also acknowledges the “Xiaomi Young Scholar Program” at Wuhan University. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).